one number is 1o more than another.the sum of twice the larger and three times the smaller is 170. what is the smaller number?

Found 2 solutions by mananth, nairb ogabar:
Answer by mananth(16946)

smaller number = x
bigger number = y
...
x+10=y
x-y=-10.....1
2y+3x=170...2
...
multiply (1) by -2
-2y-2x=20
add to (2)
5x=190
/5
x= 38
38, 48

Answer by nairb ogabar(1)

1st number.
3x
2nd number.
2(x+10)
3x + 2(x+10) = 170
3x + 2x +20 = 170
transpose 20 to the other side.
3x + 2x =170 - 20
5x = 150
divide both sides by 5
x=30